Chapter 28, Problem 003 An electron that has a velocity with component 2.5 × 106 m s and y component 3 3 x 106 m s moves through a uniform magnetic e d it x component 123 and y component a 9 T. (a) Find the magnitude of the magnetic force on the electron. (b) Repeat your calculation for a proton having the same velocity. (a) Number (b) Number Units Units
Answer #1

28] a]

The magnetic force on the electron will be:

F=qv × B

F=-e(10°)| 2.5 0.023 -0.19 0 0!=-e(10°)|一2.5(0.19)-3.3(0.023)K 3.3

=> F-+ (8.814 x 10-4N) 14 Ar

the magnitude of the magnetic force is F = 8.814 x 10-14 N

b] Since a proton and an electron have the same charge and the magnetic force is not dependent on the mass of the charge, therefore, the magnitude of charge will be the same.